Find the sum of the vectors. vector u is (4,7) and vector m (8,17). Please help.
Answer by 10021987(2)

Well, first you have to add vector u which is (4,7) and 4+7=11 and then you have to add vector m which is (8,17) and 8+17=25 then you have to add 11+25 together and then when you add them together then you will get 11+25=36 because the word sum means to add. My pleasure to help you hope the answer is correct.